) If 𝐼𝑛 = ∫ π‘π‘œπ‘ π‘›π‘₯ cos 𝑛π‘₯ 𝑑π‘₯

πœ‹

2


then find the reduction formulaΒ 

connecting 𝐼𝑛 and πΌπ‘›βˆ’1 and hence prove that 𝐼𝑛 =

πœ‹

2

𝑛+1

Β 


1
Expert's answer
2021-06-09T17:12:42-0400

Let

"I_{m}=\int\cos ^m(x)dx=\int\cos ^{m-1}(x)\cos(x)dx"

Integration by parts


"\int udv=uv-\int vdu"

"u=\cos^{m-1}( x), du=-(m-1)\cos^{m-2}(x)\cdot\sin xdx"

"dv=\cos(x)dx, v=\sin(x)"

"I_{m}=\int\cos ^m(x)dx=\int\cos ^{m-1}(x)\cos(x)dx"

"=\cos^{m-1}( x)\sin (x)+(m-1)\int\cos ^{m-2}(x)\sin^2(x)dx"

"=\cos^{m-1}( x)\sin (x)+(m-1)\int\cos ^{m-2}(x)dx"

"-(m-1)\int\cos ^{m-2}(x)\cos^2(x)dx"

"=\cos^{m-1}( x)\sin (x)+(m-1)I_{m-2}-(m-1)I_{m}, m\geq 2"

Solve for "I_{m}"


"I_{m}=\dfrac{\cos^{m-1}( x)\sin (x)}{m}+\dfrac{m-1}{m}I_{m-2}"

Theerefore

"I_n=\displaystyle\int_{0}^{\pi/2}\cos ^n(x)dx"

"=\bigg[\dfrac{\cos^{n-1}( x)\sin (x)}{n}\bigg]\begin{matrix}
 \pi/2 \\
 0
\end{matrix}+\dfrac{n-1}{n}I_{n-2}"

"=\dfrac{n-1}{n}I_{n-2}=\dfrac{n-1}{n}\displaystyle\int_{0}^{\pi/2}\cos ^{n-2}(x)dx, n\geq2"

"I_n=\displaystyle\int_{0}^{\pi/2}\cos ^n(x)dx=\dfrac{n-1}{n}\displaystyle\int_{0}^{\pi/2}\cos ^{n-2}(x)dx"

"=\dfrac{n-1}{n}I_{n-2}, n\geq 2"

"n=0: I_0=\displaystyle\int_{0}^{\pi/2}\cos ^0(x)dx=[x]\begin{matrix}
 \pi/2 \\
 0
\end{matrix}=\dfrac{\pi}{2}"

"n=1: I_0=\displaystyle\int_{0}^{\pi/2}\cos ^1(x)dx=[\sin(x)]\begin{matrix}
 \pi/2 \\
 0
\end{matrix}=1"